OpenStudy (anonymous):

Larry and Peggy are making decisions on their bank accounts. Larry wants to put the original money in an account with a higher interest rate. Peggy wants to put more money in as a principle amount because the more you start with, the more interest you will gain. Explain which method will result in more money. @inkyvoyd

OpenStudy (owlcoffee):

Oh, my knowledge in financial mathematics is very vague, I'm afraid I cannot help you :(

OpenStudy (anonymous):

do you know anybody that does?

OpenStudy (anonymous):

@phi

OpenStudy (callisto):

Is there more information provided in the question?

OpenStudy (anonymous):

nope this is the question, this chapter was about exponential and log functions.

OpenStudy (anonymous):

so thats how I would put it @douglaswinslowcooper

OpenStudy (callisto):

I think you cannot determine which method gives you higher return without knowing the exact (increased) principle, interest rate and the time of deposit.

OpenStudy (anonymous):

Yes, @Callisto, that's why we can only show them how to calculate it.

OpenStudy (callisto):

Let P be the original principle, P' be the new(more) one, r be the original interest rate, r' be the new(higher) one, t be the time of deposit. A = Pe^(rt) Larry's Amount = Pe^(r't) Peggy's Amount = P' e^(rt) When their amounts are equal, \[\frac{Pe^{r't}}{P' e^{rt}} = 1\]\[\frac{P}{P'}e^{r't-rt} = 1\]\[e^{r't-rt} = \frac{P'}{P}\]\[e^{(r'-r)t} = \frac{P'}{P}\]\[ (r'-r)t = \ln(\frac{P'}{P})\]\[ t = \frac{\ln(\frac{P'}{P})}{(r'-r)}\] When Larry's amount > Peggy's amount, \[\frac{Pe^{r't}}{P' e^{rt}} > 1\]\[\frac{P}{P'}e^{r't-rt} > 1\]\[e^{r't-rt} > \frac{P'}{P}\]\[e^{(r'-r)t} > \frac{P'}{P}\]\[ (r'-r)t > \ln(\frac{P'}{P})\]\[ t > \frac{\ln(\frac{P'}{P})}{(r'-r)}\] So, when \(t > \frac{\ln(\frac{P'}{P})}{(r'-r)}\), Larry's method will result in more $$ *Note: ln (P'/P) > 0 since P' > P ln (e^(r'-r)t) > 0 since r'-r > 0